It is necessary to have a 40% antifreeze solution in the radiator of a certian car. The radiator now has 60 liters of 20% solution. How many liters of this should be drained and replaced with 100% antifreeze to get the desired strength?
Answer by mananth(16946)

Tank contains 60 liters - 20%
let the amount to be drained be x liters
Balance will be 60-x liters --20%
100% added will be x liters
total will be 60 liters 40%
..
concentration equation
1x+0.2(60-x)=60*0.4
1x+12-0.2x=24
0.8x=12
x=12/0.8
x=15 liters will have to be drained